OpenStudy (thebadyboy):

Solve. -1/3+W= 5 3/8 Express the answer in simplest form. w =

OpenStudy (thebadyboy):

\[5\frac{ 3}{ 8}\]

OpenStudy (thebadyboy):

r u sure because im suppose to get a mixed fraction here

OpenStudy (thebadyboy):

@Allenismyname

OpenStudy (thebadyboy):

do u guys know it

OpenStudy (lasttccasey):

\[W = 5\frac{ 3 }{ 8 }+\frac{ 1 }{ 3 }\]\[ = \frac{ 43 }{ 8 }+\frac{ 1 }{ 3 }\]\[= \frac{ 129 }{ 24 }+\frac{ 8 }{ 24 }\]\[= \frac{ 137 }{ 24 }\]\[= 5\frac{ 17 }{ 24 }\]

OpenStudy (thebadyboy):

thx

OpenStudy (lasttccasey):

Do you understand how to get it?

OpenStudy (thebadyboy):

wait wouldnt it be \[5\frac{ 1}{ 24}\]

OpenStudy (lasttccasey):

Do you know how to convert a fraction to a mixed number?

OpenStudy (thebadyboy):

yes

OpenStudy (thebadyboy):

wait your right it would be 5 17/24